Providing for doubtful accounts At the end of the current year, the accounts receivable account has...

  1. Providing for doubtful accounts

    At the end of the current year, the accounts receivable account has a debit balance of $1,996,000 and sales for the year total $27,950,000. Determine the amount of the adjusting entry to provide for doubtful accounts under each of the following assumptions:

    1. The allowance account before adjustment has a debit balance of $13,600. Bad debt expense is estimated at 1/2 of 1% of sales.
    2. The allowance account before adjustment has a debit balance of $13,600. An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $29,170.
    3. The allowance account before adjustment has a credit balance of $6,200. Bad debt expense is estimated at 3/4 of 1% of sales.
    4. The allowance account before adjustment has a credit balance of $6,200. An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $51,500.

    This information has been collected in the Microsoft Excel Online file. Open the spreadsheet, perform the required analysis, and input your answers in the questions below.

    Open spreadsheet

    Round your answers to the nearest dollar.

    1. $  

    2. $  

    3. $  

    4. $
Answer #1
a 139750 =27950000*1/2*1%
b 42770 =29170+13600
c 209625 =27950000*3/4*1%
d 45300 =51500-6200